OpenStudy (anonymous):

When subtracting a proper fraction like 1/3 from 12....... Making the question look like 12-1/3........do you flip the fraction around and then divide? Making it a 3? Then the answer is 9?

OpenStudy (amistre64):

why, pray tell, would you divde in order to subtract?

OpenStudy (amistre64):

if you have 12 dollars and some takes less than a dollar away from you; would you be left with 9 dollars?

OpenStudy (anonymous):

Well the question is p=1/3 (12-p)

OpenStudy (anonymous):

Apparently the answer is 9 And I want to know why

OpenStudy (amistre64):

multiply both sides by 3 3p = 3/3 (12-p) 3p = 12-p ; add off the p +p +p ---------- 4p = 12 ; divide off the 4 p = 12/4 = 3 ... not 9

OpenStudy (anonymous):

Ty